OpenStudy (anonymous):

Jermaine has a total of 36 video games. He has 8 more car racing video games than all of the other types combined. How many car racing video games does Jermaine have? A. 28 video games <---- my answer B. 14 video games C. 26 video games D. 22 video games

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

x + x + 8 = 36 2x + 8 = 36 2x = 28 x = 14 14 + 8 =

OpenStudy (anonymous):

22

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

yep :)